OpenStudy (anonymous):

Simplify. 6a - {b - [3a - (2b + c) + 4a - (a + 2b - c)]}. 12a - 5b 6a - b + 2c 12a - 5b + 2c b

OpenStudy (kropot72):

First remove the brackets, working from the inside to the outside. Can you do that?

OpenStudy (anonymous):

ill try one sec

OpenStudy (anonymous):

um with alittle help yes

OpenStudy (kropot72):

What is the result when you remove the brackets from -(2b + c)

OpenStudy (anonymous):

ok i guess this is harder then i thought

OpenStudy (kropot72):

-(2b + c) is mathematical shorthand for( -1 * 2b) + (-1 * b) What do you get when you multiply -1 and 2b ?

OpenStudy (anonymous):

-2b

OpenStudy (kropot72):

Correct. Now multiply -1 and b.

OpenStudy (anonymous):

-b

OpenStudy (kropot72):

Right !. So -(2b + c) = -2b - c Now remove the brackets from -(a +2b - c) by multiplying each term by the -1 that is assumed outside the leading bracket.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-a-2b+c

OpenStudy (kropot72):

Absolutely right. So now we have inside the square brackets: [3a - 2b - c + 4a - a - 2b + c] Can you now collect the like terms and simplify inside the square brackets.

OpenStudy (anonymous):

ya

OpenStudy (kropot72):

What do you get when you add 3a + 4a - a

OpenStudy (anonymous):

6a

OpenStudy (kropot72):

Right. So now we have inside the square brackets [6a - 2b - 2b - c + c] So what are you left with when you add the terms in b and the terms in c inside the square brackets?

OpenStudy (anonymous):

b cansals right

OpenStudy (kropot72):

Yes. b disappears from inside the square brackets.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6a-2c

OpenStudy (kropot72):

Sorry. I should have said that c cancels. + c - c = 0 -2b -2b = -4b So inside the curly brackets we have: {b - [6a - 4b]} = {b - 6a + 4b} Do you follow so far?

OpenStudy (anonymous):

yup

OpenStudy (kropot72):

We're nearly there! So putting all the terms that are left together we have: 6a - {b - 6a + 4b} = ? when you multiply everything inside the curly brackets by -1 to enable the curly brackets to be removed. Can you do this?

OpenStudy (anonymous):

-b+6a-4b

OpenStudy (kropot72):

Right again. You're getting good at this! So finally we have: 6a - b + 6a -4b = ? You add the terms in a and add the terms in b (observing their signs) and you should have the correct answer.

OpenStudy (anonymous):

12a-5b

OpenStudy (kropot72):

Well done :) That is one of the choices.

OpenStudy (anonymous):

ok thanks

OpenStudy (kropot72):

You're welcome :)